OpenStudy (anonymous):

Will Fan! Equation C: y = 2x + 8 Equation D: y = 2x + 2 Which of the following best describes the solution to the given set of equations? No solution One solution Two solutions Infinite solutions

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

\[2x + 2= 2x + 8\] Solve this

OpenStudy (anonymous):

i cant

OpenStudy (kidrah69):

Yea you can.

OpenStudy (anonymous):

dont i have to bring the 2x over?

OpenStudy (anonymous):

when i do that i get 0

OpenStudy (anonymous):

and then im stuck

OpenStudy (kidrah69):

0=0?

OpenStudy (anonymous):

here's what i do 2x + 2 = 2x + 8 -2 -2 2x = 2x +6 and now im stuck

OpenStudy (kidrah69):

2x+2x=?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

no wait its 4 but that snot how you do it. is it? i thought you subtract 2x by 2x

OpenStudy (anonymous):

* thats not

OpenStudy (kidrah69):

4x=0 4/0= ?

OpenStudy (anonymous):

what?

OpenStudy (kidrah69):

2+2=4 which makes 4x=6 4/6 is the answer x=4/6 so what is the FINAL answer? How many solutions?

OpenStudy (anonymous):

there's one solution.

OpenStudy (kidrah69):

Good thats the answer

OpenStudy (anonymous):

thanks! i have another question can you help?

OpenStudy (kidrah69):

Sure. :)

OpenStudy (anonymous):

okay im gonna make a new post